OpenStudy (anonymous):

Given segment AM as the perpendicular bisector in the equilateral triangle ABC, which statement is true? BM + CM = 2AM AB + BC = 4AM 2BM + 2CM = 2AB 2BM + 3CM = 5AB .

OpenStudy (amistre64):

bm + cm = ab; am = sqrt(3)/2; bm=1; cm=1; i wonder

OpenStudy (blacksteel):

Since AM is the perpendicular bisector of the equilateral triangle, M should divide the line BC in half. Thus BM = CM, and BM + CM = BC. Thus, 2BM + 2CM = 2(BM + CM) = 2BC, and since the triangle is equilateral, 2BC = 2AB. Thus, the third option is correct.

OpenStudy (amistre64):

ab would equal say 2; ab + bc = 4; sqrt(3/2)*4 not the same

OpenStudy (anonymous):

i believe its the 3rd one. 2(BM + CM) = 2(AB)

OpenStudy (anonymous):

thank you :)

myininaya (myininaya):

i agree with the third one since we have 2BM+2MC=2*x/2+2*x/2=x+x=2x=2(AB)

myininaya (myininaya):

where x is length of BC and AB and AC